OpenStudy (anonymous):

write in standard form. x^2+7x+12=0

OpenStudy (anonymous):

You can either use the quadratic formula or the splitting method, whichever you find easier. By the splitting method: The given equation can be expressed as: \[x ^{2} + 4x + 3x + 12 = 0\] \[x( x+4) + 3(x+4) = 0\] \[(x+3) (x+4) = 0\] Therefore, x = -4 or x = -3